Consider the following project for Dawg Incorporated:

YEAR

0

1

2

3

4

5

Sales

$150,000

$150,000

$150,000

$150,000

$150,000

Cost of Goods

$70,000

$70,000

$70,000

$70,000

$70,000

S&A

$30,000

$30,000

$30,000

$30,000

$30,000

Depreciation

$30,000

$30,000

$30,000

$30,000

$30,000

Investment in NWC

$1,000

$500

$500

$500

$500

$500

Investment in Gross PPE

$150,000

The project will last 5 years and has the same risk as the typical Dawg Incorporated project.  The firm has a capital structure of 40.00% debt and 60.00% equity. The cost of debt is 8.00%, while the cost of equity is estimated at 15.00%. The tax rate facing the firm is 30.00%.  There is no reclaimed NWC or NSV at end of year 5.

What is the NPV for the project?
Answer #1

In order to calculate NPV first we have to find Weighted Average Cost of Capital

Cost of Equity = 15%

Cost of Debt = Interest *(1-Tax) = 8%*(1-30%) = 5.6%

Sources Weights (A=X/Y) Cost (B) WACC (A*B)
Debt 0.40 5.6% 0.0224
Equity 0.60 15% 0.0900
Total (Y) WACC (SUM OF (A*B)) 0.1124

In first case i am taking Debt as in 40% 150,000$, hence interest would be 8% of it = 150000*40%*8% = 4,800$

Present Vale Factor = 1/(1+WACC)^Number of years to be bought back

Calculation of Net Present Value (In Dollars)
Particulars 0 1 2 3 4 5
Sales 150000 150000 150000 150000 150000
Less: Cost of Goods Sold -70000 -70000 -70000 -70000 -70000
Less: Selling and Administration -30000 -30000 -30000 -30000 -30000
Earnings Before Interest Tax and Depriciation 50000 50000 50000 50000 50000
Less: Depreciation -30000 -30000 -30000 -30000 -30000
Earnings Before Interest and Tax 20000 20000 20000 20000 20000
Less: Interest (150000*40%*8%) -4800 -4800 -4800 -4800 -4800
Earnings Before Tax 15200 15200 15200 15200 15200
Less: Tax Rate @30% -4560 -4560 -4560 -4560 -4560
Profit After Tax 10640 10640 10640 10640 10640
Add: Interest *(1-Tax) 3360 3360 3360 3360 3360
Unlevered Income 14000 14000 14000 14000 14000
Add: Depreciation 30000 30000 30000 30000 30000
Net Operating Profit After Tax 44000 44000 44000 44000 44000
Less: Capital Investment -150000
Less: Net Working Capital -1000 -500 -500 -500 -500 -500
Free Cash Flow (A) -151000 43500 43500 43500 43500 43500
Present Value Factor @ 11.24% (B) 1 0.89895721 0.80812406 0.726468954 0.653065 0.587077
Present Values (A*B) -151000 39104.63862 35153.3968 31601.39951 28408.31 25537.85
Net Present Value (Sum of (A*B)) 8805.592299

In Second case i am taking Debt as in 40% 151,000$(Capital Investment + Working Capital), hence interest would be 8% of it = 151000*40%*8% = 4,832$

Calculation of Net Present Value (In Dollars)
Particulars 0 1 2 3 4 5
Sales 150000 150000 150000 150000 150000
Less: Cost of Goods Sold -70000 -70000 -70000 -70000 -70000
Less: Selling and Administration -30000 -30000 -30000 -30000 -30000
Earnings Before Interest Tax and Depriciation 50000 50000 50000 50000 50000
Less: Depreciation -30000 -30000 -30000 -30000 -30000
Earnings Before Interest and Tax 20000 20000 20000 20000 20000
Less: Interest (150000*40%*8%) -4832 -4832 -4832 -4832 -4832
Earnings Before Tax 15168 15168 15168 15168 15168
Less: Tax Rate @30% -4550.4 -4550.4 -4550.4 -4550.4 -4550.4
Profit After Tax 10617.6 10617.6 10617.6 10617.6 10617.6
Add: Interest *(1-Tax) 3382.4 3382.4 3382.4 3382.4 3382.4
Unlevered Income 14000 14000 14000 14000 14000
Add: Depreciation 30000 30000 30000 30000 30000
Net Operating Profit After Tax 44000 44000 44000 44000 44000
Less: Capital Investment -150000
Less: Net Working Capital -1000 -500 -500 -500 -500 -500
Free Cash Flow (A) -151000 43500 43500 43500 43500 43500
Present Value Factor @ 11.24% (B)